OpenStudy (anonymous):

Find the exact value of tan (arcsin (2/5)).

OpenStudy (anonymous):

0.4364

OpenStudy (anonymous):

@AFleming42 got it..??

OpenStudy (anonymous):

what's the answer in terms of pi? @Kashan

OpenStudy (hba):

let [arcsin(2/5)] = y hence: siny = 2/5 now, rewriting tany in terms of siny (whose value is known), you get: tany = siny/cosy = siny/[±√(1 - sin²y)] owing to arcsine function range (-pi/2,+pi/2), being siny positive, y belongs to the 1st quadrant, thus cosy is positive too; therefore, taking the plus sign, tany = siny/√(1 - sin²y) = (2/5)/√[1 - (2/5)²]

OpenStudy (hba):

= (2/5)/√[1 - (4/25)]

OpenStudy (hba):

=(2/5)/√[(25-4)/25]

OpenStudy (hba):

=(2/5)/√(21/25)

OpenStudy (anonymous):

so complicated :o

OpenStudy (hba):

Well It Is Not At All Complicated :(

OpenStudy (hba):

Arcsin(2/5) = arcsin(y/r) ⇒ x = √(5² - 2²) = √(25 - 4) = √21 ⇒ tan(arcsin(2/5)) = y/x = 2/√21